OpenStudy (anonymous):

find the perimeter of a right triangle with legs that measure 5 cm and 9 cm, round to the nearest tenth.

OpenStudy (jiteshmeghwal9):

perimeter = side + side + side

OpenStudy (jiteshmeghwal9):

now find the unknown side using pythagoras theorem

OpenStudy (jiteshmeghwal9):

sqrt{5^2+9^2}=unknown side

OpenStudy (jiteshmeghwal9):

sqrt{25+81} sqrt{106}

OpenStudy (jiteshmeghwal9):

first of all tell me is there hypotenuse unknown

OpenStudy (jiteshmeghwal9):

???

OpenStudy (anonymous):

i typed the only information given to me.

OpenStudy (lgbasallote):

you're actually supposed to solve for the hypotenuse

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

what formula do i use?

OpenStudy (anonymous):

A62=B62=C^2 5^2+9^2=C^2 25+81=C^2 C^2=106 C=SQRT106 C=10.3 5+9+10.3=24.3 ANS. FOR THE PERIMETER. is this right?

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

thank you